Circular No. 8661

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006H, 2006I, 2006J, 2006K, 2006L, 2006M, 2006N Six apparent supernovae have been found on unfiltered CCD images:

2006I by T. Puckett and T. Crowley (cf. IAUC 8660), 2006J by N. J. Ponticello and W. Li (LOSS/KAIT; cf. IAUC 8660), 2006N by M. Armstrong (cf. IAUC 8643), and the rest by the “Nearby Supernova Factory” (NSF) collaboration (via NEAT images, Palomar 1.2-m telescope; cf. CBET 374).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006I Jan. 19.42 10^{h}59^{m}30.^{s}03 +75^{o}11′34′′.0 18.1 5′′.1 W, 4′′.9 N
2006J Jan. 20.09 0 20 20.15 + 7 42 07.0 18.3 0′′.5 E, 4′′.7 N
2006K Jan. 17.5 11 29 46.26 −17 25 17.8 18.6  —
2006L Jan. 17.6 13 07 11.76 +28 38 46.2 19.9  —
2006M Jan. 17.6 13 27 19.76 +31 47 14.5 18.5  —
2006N Jan. 21.02 6 08 31.24 +64 43 25.1 15.0 0′′.4 E, 0′′.1 N

Additional magnitudes of 2006I in NGC 3465: Jan. 8.46 UT, [19.2; 20.42, 18.0. SN 2004bc also appeared in NGC 3465 (IAUC 8315). Additional KAIT magnitudes of 2006J in IC 13: 2005 Oct. 29.15, [19.5; Dec. 17.17, [18.5; 2006 Jan. 21.09, 18.3. Additional magnitudes of 2006N in MCG +11-8-12: 2004 Aug. 26, [19.5; 21.738, 15.0. Nothing was visible at the location of 2006N on Palomar Sky Survey images from 1996 (limiting red mag 20.8) and 1993 (limiting blue mag 22.5).

NSF spectroscopy (range 320–1000 nm, University of Hawaii 2.2-m telescope; cf. CBET 374) of 2006M in PGC 47137 and 2006L show them to be type-IIn supernovae, while 2006K is a type-II supernova. NSF spectroscopy of 2006H (IAUC 8660), obtained on Jan. 20.3 UT, shows it to be a type-Ia supernova; its spectrum was similar to the peculiar, subluminous SN 1991bg (Filippenko et al. 1992, A.J. 104, 1543) at two days past maximum light (cf. CBET 371).

S. Blondin, M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum (range 350–740 nm) of SN 2006N, obtained on Jan. 22.23 UT by P. Berlind with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-Ia supernova, with a spectral-feature age (Riess et al. 1997, A.J. 114, 722) of −3 days from maximum light. Adopting a recession velocity of 4280 km/s for the host galaxy (Falco et al. 1999, PASP 111, 438), the maximum absorption in the Si II line (rest 635.5 nm) is blueshifted by 12000 km/s.

2006 January 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green